ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 32, 41 Abe, Yukitaka, Toyama University, Japan Universal functions in several variables An entire function whose compositions with translations are dense is said to be universal. G. D. Birkhoff, in 1929, was the first to prove the existence of universal functions on the complex plane. Since then several universalities were defined and studied in one variable. We consider universal functions in several variables. We extend several results in one variable to the multi-dimensional case. Part of these results is joint work with P. Zappa. References: Y. Abe, Universal holomorphic functions in several variables, Analysis 17(1997), 71-77. Y. Abe and P. Zappa, Universal functions on complex general linear groups, preprint.